# On some indices of foliations and applications

### v1

@@ Abstract.
In this paper we establish a relationship between the Milnor number, the $\chi$-number, and the Tjurina number of a foliation with respect to an effective balanced divisor of separatrices. Moreover, using the Gómez-Mont–Seade–Verjovsky index, we prove that the difference between the multiplicity and the Tjurina number of a foliation with respect to a reduced curve is independent of the foliation. We also derive an adjunction formula for the Tjurina number of a foliation with respect to a reduced curve. These local results have global consequences; for instance, we provide a new proof of a global result regarding the multiplicity of a foliation due to Cerveau-Lins Neto and a new proof of a Soares’s inequality for the sum of the Milnor number of an invariant curve of a foliation. Additionally, we obtain upper and lower bounds for the global Tjurina number of a foliation on the complex projective plane. Finally, we provide an example that answers a conjecture posed by Alcántara–Mozo-Fernández about foliations on the complex projective plane having a unique singularity.
@@
Holomorphic foliations, Milnor number, Tjurina number, Multiplicity of a foliation along a divisor of separatrices, Dimca-Greuel bound
@@
The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done (Spanish grants PID2019-105896GB-I00 and PID2023-149508NB-I00 funded). This work was also funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2024-PI1100. The first author acknowledges support from CNPq Projeto Universal 408687/2023-1 ”Geometria das Equações Diferenciais Algébricas” and CNPq-Brazil PQ-306011/2023-9. We are grateful to Marcelo E. Hernandes and Hernán Neciosup Puican for their remarks.
@@ 1. Introduction
First, using known adjunction formulas for the GSV-index and the Tjurina number of a singular curve, we present an adjunction formula for the Tjurina number of a foliation, see for instance Proposition 3.1. Next, we prove in Proposition 4.1 that the difference between the multiplicity and the Tjurina number of a foliation with respect to a reduced curve does not depend on the foliation. More specifically, let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$ and let $C$ be any $\mathcal{F}$-invariant reduced curve. Then
@@
where $\mu_{p}(\mathcal{F},C)$ and $\tau_{p}(\mathcal{F},C)$ denote the multiplicity and Tjurina number of $\mathcal{F}$ along $C$, respectively, and $\mu_{p}(C)$, $\tau_{p}(C)$ are the classical Milnor and Tjurina number of $C$. This result leads to several applications; for instance, we show that
As a consequence in Corollary 4.2, we provide a positive answer to a question posed in [17].
In Section 5, following the terminology of balaced divisor of separatrices of a foliation introduced by Genzmer [21], we present Proposition 5.1 that generalizes Corollary B of [16]. We recall that a foliation $\mathcal{F}$ is said to be a generalized curve if there are no saddle-nodes in its reduction process of singularities. This concept was introduced by Camacho, Lins Neto and Sad in [9] and the non-dicritical case, delimits a family of foliations whose topology is closely related to that of their separatrices (i.e. local invariant curves). Generalized curve foliations are part of the broader family of second type foliations, introduced by Mattei and Salem in [30]. Foliations in this family may admit saddle-nodes in the reduction process of singularities, provided that they are not tangent saddle-nodes (see [18, Definition 2.1]). In order to study second type foliations, the $\chi$-number of $\mathcal{F}$ was introduced in [16, Section 3]. This number, denoted by $\chi_{p}(\mathcal{F})$, can be interpreted as the difference $\mu_{p}(\mathcal{F})-\mu_{p}(\mathcal{F},\mathcal{B})$, where $\mu_{p}(\mathcal{F})$ is the Milnor number of $\mathcal{F}$ at $p$ and $\mu_{p}(\mathcal{F},\mathcal{B})$ is the multiplicity of $\mathcal{F}$ along a balanced divisor of separatrices $\mathcal{B}$.
@@
The proof of Theorem A will be given in Section 6. Finally, we apply our local results in three ways. First, we provide a new proof of a result by Cerveau-Lins Neto [11, Proposition 7.1] and a new proof of a Soares’s inequality [35, Theorem 7.3] for the sum of the Milnor number of an invariant curve of a foliation. In Theorem 7.5, we establish upper and lower bounds for the global Tjurina number of a foliation on the complex projective plane. Lastly, under certain conditions, we answer a question posed by Alcántara–Mozo-Fernández [1, p. 16] regarding the existence of non-dicritical logarithmic foliations with a unique singular point and an invariant algebraic curve passing through the singularity.
@@ 2. Basic tools
To establish the terminology and the notation, we recall some basic concepts from local Foliation Theory. Unless stated otherwise, throughout this text, $\mathcal{F}$ denotes a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. In local coordinates $(x,y)$ centered at $p$, the foliation $\mathcal{F}$ is given by a holomorphic 1-form
@@
for some $h\in\mathbb{C}[[x,y]]$. If $C$ is an irreducible $\mathcal{F}$-invariant curve then we will say that $C$ is a separatrix of $\mathcal{F}$. The separatrix $C$ is analytical if $f$ is convergent. We denote by $\mathrm{Sep}_{p}(\mathcal{F})$ the set of all separatrices of $\mathcal{F}$. When $\mathrm{Sep}_{p}(\mathcal{F})$ is a finite set we will say that the foliation $\mathcal{F}$ is non-dicritical and we call total union of separatrices of $\mathcal{F}$ to the union of all elements of $\mathrm{Sep}_{p}(\mathcal{F})$. Otherwise we will say that $\mathcal{F}$ is a dicritical foliation.
@@ 2.1. GSV-index
The GSV-index was introduced by Gómez-Mont-Seade-Verjovsky in [25]. This index is a topological invariant for vector fields on singular complex varieties and is closely related to the Euler-Poincaré characteristic of the Milnor fiber (see, for instance [8, Theorem 3.1]). It also relates to the Schwartz index and local Euler obstruction through a proportionality theorem [5, Theorem 3.1]. In this subsection, we follow the presentation of [6].
Let $\mathcal{F}:\omega=0$ be a singular foliation at $(\mathbb{C}^{2},p)$. Let $C:f(x,y)=0$ be an $\mathcal{F}$-invariant curve, where $f(x,y)\in\mathbb{C}[[x,y]]$ is reduced. Then, as in the convergent case, there are $g,h\in\mathbb{C}[[x,y]]$ (depending on $f$ and $\omega$), with $f$ and $g$ and $f$ and $h$ relatively prime and a $1$-form $\eta$ (see [36, Lemma 1.1 and its proof]) such that
@@
The Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ (GSV-index) with respect to an analytic $\mathcal{F}$-invariant curve $C$ is
@@
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$. Camacho-Lins Neto-Sad [9, Section 4] defined the multiplicity of $\mathcal{F}$ along $B$ at $p$ as $\mu_{p}(\mathcal{F},B):=\text{ord}_{t}\theta(t)$, where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, we have
@@ 2.3. The Tjurina number of a foliation along an invariant reduced curve
This number appears for first time in [24], and the terminology of Tjurina number of $\mathcal{F}$ with respect to $C$ was given in [10, p. 159].
3. Adjunction formulas for the Tjurina number of a foliation
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$, and let $C=\cup_{j=1}^{r}C_{j}$ be an $\mathcal{F}$-invariant reduced curve such that each $C_{j}$ is irreducible. Then, from [38, Theorem 6.5.1], [27, Proposition 3.2], [29, Equation (2.2), p. 329], and [7, p. 29], we obtain the following formulas, respectively:
@@ 3. Adjunction formulas for the Tjurina number of a foliation
where $I_{p}(C_{i},C_{j})$ denotes the intersection multiplicity of the branches $C_{i}$ and $C_{j}$ at $p$, and $\Theta$ is a number that depends on the module of Kähler differentials of $C$. Its explicity formula is given in [27, Theorem 5.1].
Analogous to the adjunction formula for the Tjurina number of curves, we provide an adjunction formula for the Tjurina number of a foliation $\mathcal{F}$ with respect to a reduced $\mathcal{F}$-invariant curve $C$, in terms of the Tjurina numbers of $\mathcal{F}$ with respect to the irreducible components of $C$.
@@
$\tau_{p}(\mathcal{F},C)$ $=$ $\sum^{r}_{j=1}\tau(\mathcal{F},C_{j})-\left(\sum_{1\leq i<j\leq r}I_{p}(C_{i},C_{j})\right)+\Theta.$
@@
After [16, Proposition 6.2] we get $\tau_{p}(\mathcal{F},C)=GSV_{p}(\mathcal{F},C)+\tau_{p}(C)$. Hence by equalities (6) and (4), we obtain
@@
$=$ $\sum^{r}_{j=1}\tau(\mathcal{F},C_{j})-\left(\sum_{i<j}I_{p}(C_{i},C_{j})\right)+\Theta.$
@@
We borrow from [16, Example 6.5] the family of foliations $\mathcal{F}_{k}$, $k\geq 3$, given by the $1$-form
@@
which is a family of dicritical foliations which are not of second type. An effective balanced divisor of separatrices of $\mathcal{F}_{k}$ is $\mathcal{B}=\mathcal{B}_{0}:xy=0$. Put $B_{1}:x=0$ and $B_{2}:y=0$. Hence $\tau_{0}(\mathcal{F}_{k},\mathcal{B})=3k-2$, $\tau_{0}(\mathcal{F}_{k},B_{1})=k$, and $\tau_{0}(\mathcal{F}_{k},B_{2})=2k-1$, since $\Theta=0$ by (4), we get
@@ 4. The GSV-index, the multiplicity and the Tjurina number of a foliation
Let $C=\cup_{j=1}^{r}C_{j}$. It follows from [17, Eq. 5] that
@@
By combining (9), (3), (5), and (6), we obtain $GSV_{p}(\mathcal{F},C)=\mu_{p}(\mathcal{F},C)-\mu(C)$. Finally, using [16, Proposition 6.2], we get
@@
After Proposition 4.1 and [2, Theorem 3.2], we provide a positive answer to Question 1 posed in [17].
@@
(10) $\mu_{p}(\mathcal{F},C)-\tau_{p}(\mathcal{F},C)=\mu_{p}(C)-\tau_{p}(C)<\mu_{p}(C)/4.$
By (7), the proof follows substituing into (10). ∎
@@
The corollary follows from Proposition 4.1, and [33, Satz p.123, (a) and (d)]. ∎
@@
Applying Proposition 4.1 to each $C_{i}$ and to $C$, and substituing into (11), we obtain the desired inequality. ∎
@@ 4.1. Values set of fractional ideals and the Tjurina number of a foliation
In this subsection, we follow the notations from [3]. Let $C:f(x,y)=0$ be a (reduced) germ of complex plane curve singularity at $p\in\mathbb{C}^{2}$ with equation $f=\Pi_{i=1}^{r}f_{i}=0$, where $f_{i}\in\mathbb{C}\{x,y\}$ is irreducible. Each $f_{i}$ defines a branch $C_{i}$ and its analytic type is characterized by the local ring $\mathcal{O}_{i}:=\mathbb{C}\{x,y\}/(f_{i})$ up to $\mathbb{C}$-algebra isomorphism. The fractions field $\mathcal{H}_{i}$ of $\mathcal{O}_{i}$ is isomorphic to $\mathbb{C}(t_{i})$ and associated to it we have a canonical discrete valuation $\nu_{i}:\mathcal{H}_{i}\to\overline{\mathbb{Z}}:=\mathbb{Z}\cup\{\infty\}$. If $\overline{h}\in\mathcal{O}_{i}$ with $h\in\mathbb{C}\{x,y\}$, then we have $\nu_{i}(\overline{h})=I_{p}(f_{i},h)$. The image of $\nu_{i}$ of $\mathcal{O}_{i}\setminus\{0\}$, that is
@@
is the semigroup of values of $C_{i}$. Put $\mathcal{O}:=\mathbb{C}\{x,y\}/(f)$. The semigroup of values of $C$ is
@@
Let $C$ be a germ of reduced plane curve at $p\in\mathbb{C}^{2}$. With the previous notation, put $\overline{\varLambda}=\varLambda_{f}\cup\{0\}$. Then, we have

### v2

@@ Abstract.
In this paper we establish a relationship between the Milnor number, the $\chi$-number, and the Tjurina number of a foliation with respect to an effective balanced divisor of separatrices. Moreover, using the Gómez-Mont–Seade–Verjovsky index, we prove that the difference between the multiplicity and the Tjurina number of a foliation with respect to a reduced curve is independent of the foliation. We also derive a local formula for the Tjurina number of a foliation with respect to a reduced curve. From a global point of view, these results lead to the following consequences: we provide a new proof of a global result regarding the multiplicity of a foliation due to Cerveau-Lins Neto and a new proof of a Soares’s inequality for the sum of the Milnor number of an invariant curve of a foliation. Additionally, we obtain bounds for the global Tjurina number of a foliation on the complex projective plane. Finally, we provide an answer to the conjecture posed by Alcántara and Mozo-Fernández about foliations on the complex projective plane having a unique singularity.
@@
Holomorphic foliations, Milnor number, Tjurina number, Multiplicity of a foliation along a divisor of separatrices, GSV-index.
@@ 1. Introduction
First, using known local formulas for the GSV-index and the Tjurina number of a singular curve, we present a local formula for the Tjurina number of a foliation (see Proposition 3.1). Next, we prove in Proposition 4.1 that the difference between the multiplicity and the Tjurina number of a foliation with respect to a reduced curve does not depend on the foliation. More specifically, let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$ and let $C$ be any $\mathcal{F}$-invariant reduced curve. Denoting by $\mu_{p}(\mathcal{F},C)$ and $\tau_{p}(\mathcal{F},C)$ the multiplicity and Tjurina number of $\mathcal{F}$ along $C$, respectively, we obtain
@@
where $\mu_{p}(C)$, $\tau_{p}(C)$ are the Milnor and Tjurina numbers of $C$. This result leads to several applications; for instance, we show that
$GSV_{p}(\mathcal{F},C)<4\tau_{p}(\mathcal{F},C)-3\mu_{p}(\mathcal{F},C),$
where $GSV_{p}(\mathcal{F},C)$ denotes the $GSV$-index of $\mathcal{F}$ along $C$ at $p$. Some applications also are presented in Sections 4 and 7, for example, we provide a positive answer to a question posed in [20].
In Section 5, following the terminology of balanced divisor of separatrices of a foliation introduced by Genzmer [24], we present Proposition 5.1 that generalizes Corollary B of [19]. We recall that a foliation $\mathcal{F}$ is said to be a generalized curve if there are no saddle-nodes in its reduction process of singularities. This concept was introduced by Camacho, Lins Neto and Sad in [9] and the non-dicritical case, delimits a family of foliations whose topology is closely related to that of their separatrices (i.e. local invariant curves). Generalized curve foliations are part of the broader family of second type foliations, introduced by Mattei and Salem in [33]. Foliations in this family may admit saddle-nodes in the reduction process of singularities, provided that they are not tangent saddle-nodes (see [21, Definition 2.1]). In order to study second type foliations, the $\chi$-number of $\mathcal{F}$ was introduced in [19, Section 3]. This number, denoted by $\chi_{p}(\mathcal{F})$, can be interpreted as the difference $\mu_{p}(\mathcal{F})-\mu_{p}(\mathcal{F},\mathcal{B})$, where $\mu_{p}(\mathcal{F})$ is the Milnor number of $\mathcal{F}$ at $p$ and $\mu_{p}(\mathcal{F},\mathcal{B})$ is the multiplicity of $\mathcal{F}$ along a balanced divisor of separatrices $\mathcal{B}$. For more details on balanced divisor of separatrices of a foliation, we refer the reader to [24] and [26].
@@
The proof of Theorem A will be given in Section 6. Finally, we apply our local results in three ways. First, we provide a new proof of a result by Cerveau-Lins Neto [12, Proposition 7.1] and a new proof of a Soares’s inequality [39, Theorem 7.3] for the sum of the Milnor number of an invariant curve of a foliation. In Theorem 7.5, we establish bounds for the global Tjurina number of a foliation on the complex projective plane. Lastly, we answer a question posed by Alcántara and Mozo-Fernández [1, p. 16] regarding the existence of non-dicritical logarithmic foliations with a unique singular point and an invariant algebraic curve passing through the singularity.
@@ 2. Basic tools
To establish the terminology and notation, we recall some basic concepts of the theory of holomorphic foliations. Unless stated otherwise, throughout this text, $\mathcal{F}$ denotes a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. In local coordinates $(x,y)$ centered at $p$, the foliation $\mathcal{F}$ is given by a holomorphic 1-form
@@
for some $h\in\mathbb{C}[[x,y]]$. If $C$ is an irreducible $\mathcal{F}$-invariant curve then we will say that $C$ is a separatrix of $\mathcal{F}$ at $p$. The separatrix $C$ is analytical if $f$ is convergent. We denote by $\mathrm{Sep}_{p}(\mathcal{F})$ the set of all separatrices of $\mathcal{F}$ at $p$. When $\mathrm{Sep}_{p}(\mathcal{F})$ is a finite set we will say that the foliation $\mathcal{F}$ is non-dicritical and the union of all elements of $\mathrm{Sep}_{p}(\mathcal{F})$ is called total union of separatrices of $\mathcal{F}$ at $p$. Otherwise, we will say that $\mathcal{F}$ is a dicritical foliation.
@@ 2.1. GSV-index
The GSV-index was introduced by Gómez-Mont-Seade-Verjovsky in [28]. This index is a topological invariant for vector fields on singular complex varieties and is closely related to the Euler-Poincaré characteristic of the Milnor fiber (see, for instance, [8, Theorem 3.1]). It also relates to the Schwartz index and local Euler obstruction through a proportionality theorem [5, Theorem 3.1]. Below we will present Brunella’s definition of the GSV-index for one dimensional holomorphic foliations (see [6]).
Let $\mathcal{F}:\omega=0$ be a singular foliation at $(\mathbb{C}^{2},p)$. Let $C:f(x,y)=0$ be an $\mathcal{F}$-invariant curve, where $f(x,y)\in\mathbb{C}[[x,y]]$ is reduced. Then, as in the convergent case, there are $g,h\in\mathbb{C}[[x,y]]$ (depending on $f$ and $\omega$), with $f$ and $g$ and $f$ and $h$ relatively prime and a $1$-form $\eta$ (see [40, Lemma 1.1 and its proof]) such that
@@
Expression (1) was first proved by K. Saito [37]. Subsequently, A. Lins Neto [32] obtained the same result in the case where $f$ is irreducible, and T. Suwa [40], in the case where $f$ is reduced, showed that the functions $f,g,h$ can be chosen in such a way that each pair $(f,g)$ and $(f,h)$ is relatively prime.
The GSV-index of the foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ with respect to an analytic $\mathcal{F}$-invariant curve $C$ is
@@
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$ at $p$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$. Camacho-Lins Neto-Sad [9, Section 4] defined the multiplicity of $\mathcal{F}$ along $B$ at $p$ as $\mu_{p}(\mathcal{F},B):=\text{ord}_{t}\theta(t)$, where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, we have
@@ 2.3. The Tjurina number of a foliation along an invariant reduced curve
This number appears for the first time in [27], and the terminology of Tjurina number of $\mathcal{F}$ with respect to $C$ was given in [10, p. 159].
3. Local formulas for the Tjurina number of a foliation
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$, and let $C=\cup_{j=1}^{r}C_{j}$ be an $\mathcal{F}$-invariant reduced curve such that each $C_{j}$ is irreducible. Then, from [42, Theorem 6.5.1], [30, Proposition 3.2], [31, Equation (2.2), p. 329], and [7, p. 29], we obtain the following formulas, respectively:
@@ 3. Local formulas for the Tjurina number of a foliation
where $I_{p}(C_{i},C_{j})$ denotes the intersection multiplicity of the branches $C_{i}$ and $C_{j}$ at $p$, and $\Theta$ is a number that depends on the module of Kähler differentials of $C$, whose explicit formula can be found in [30, Theorem 5.1]. Moreover, the Milnor number $\mu_{p}(\mathcal{F})$ of the foliation $\mathcal{F}$ at a point $p$, defined by the $1$-form $\omega=P(x,y)dx+Q(x,y)dy$, is given by $\mu_{p}(\mathcal{F})=I_{p}(P,Q).$ Recall that we assume $P$ and $Q$ to be coprime, so $\mu_{p}(\mathcal{F})$ is a non-negative integer. In [9, Theorem A], it was proved that the Milnor number of a foliation is a topological invariant.
Analogous to the local formula for the Tjurina number of curves, we provide an local formula for the Tjurina number of a foliation $\mathcal{F}$ with respect to a reduced $\mathcal{F}$-invariant curve $C$, in terms of the Tjurina numbers of $\mathcal{F}$ with respect to the irreducible components of $C$.
@@
$\tau_{p}(\mathcal{F},C)$ $=$ $\sum^{r}_{j=1}\tau_{p}(\mathcal{F},C_{j})-\left(\sum_{1\leq i<j\leq r}I_{p}(C_{i},C_{j})\right)+\Theta.$
@@
After [19, Proposition 6.2] we get $\tau_{p}(\mathcal{F},C)=GSV_{p}(\mathcal{F},C)+\tau_{p}(C)$. Hence by equalities (6) and (4), we obtain
@@
$=$ $\sum^{r}_{j=1}\tau_{p}(\mathcal{F},C_{j})-\left(\sum_{i<j}I_{p}(C_{i},C_{j})\right)+\Theta.$
@@
We borrow from [19, Example 6.5] the family of foliations $\mathcal{F}_{k}$, $k\geq 3$, given by the $1$-form
@@
which is a family of dicritical foliations which are not of second type. An effective balanced divisor of separatrices of $\mathcal{F}_{k}$ is $\mathcal{B}=\mathcal{B}_{0}:xy=0$. Let $B_{1}:x=0$ and $B_{2}:y=0$. Hence $\tau_{0}(\mathcal{F}_{k},\mathcal{B})=3k-2$, $\tau_{0}(\mathcal{F}_{k},B_{1})=k$, and $\tau_{0}(\mathcal{F}_{k},B_{2})=2k-1$, since $\Theta=0$ by (4), we get
@@ 4. The GSV-index, the multiplicity and the Tjurina number of a foliation
Let $C=\cup_{j=1}^{r}C_{j}$. It follows from [20, Eq. 5] that
@@
By combining (9), (3), (5), and (6), we obtain $GSV_{p}(\mathcal{F},C)=\mu_{p}(\mathcal{F},C)-\mu_{p}(C)$. Finally, using [19, Proposition 6.2], we get
@@
After Proposition 4.1 and [2, Theorem 3.2], we prove Corollary 4.2 which provides a positive answer to Question 1 posed in [20].
@@
$\mu_{p}(\mathcal{F},C)-\tau_{p}(\mathcal{F},C)=\mu_{p}(C)-\tau_{p}(C)<\mu_{p}(C)/4.$
and proof follows from formula (7). ∎
@@
It follows from Proposition 4.1, and [36, Satz p.123, (a) and (d)]. ∎
@@
Applying Proposition 4.1 to each $C_{i}$ and to $C$, and substituting into (10), we obtain the desired inequality. ∎
@@ 4.1. Values set of fractional ideals and the Tjurina number of a foliation
In this subsection, we follow the notations from [3]. Let $C:f(x,y)=0$ be a (reduced) germ of complex plane curve singularity at $p\in\mathbb{C}^{2}$ with equation $f=\Pi_{i=1}^{r}f_{i}=0$, where $f_{i}\in\mathbb{C}\{x,y\}$ is irreducible. Each $f_{i}$ defines a branch $C_{i}$ and its analytic type is characterized by the local ring $\mathcal{O}_{i}:=\mathbb{C}\{x,y\}/(f_{i})$ up to $\mathbb{C}$-algebra isomorphism. The field of fractions $\mathcal{H}_{i}$ of $\mathcal{O}_{i}$ is isomorphic to $\mathbb{C}(t_{i})$ and associated to it we have a canonical discrete valuation $\nu_{i}:\mathcal{H}_{i}\to\overline{\mathbb{Z}}:=\mathbb{Z}\cup\{\infty\}$. If $\overline{h}\in\mathcal{O}_{i}$ with $h\in\mathbb{C}\{x,y\}$, then we have $\nu_{i}(\overline{h})=I_{p}(f_{i},h)$. The image of $\nu_{i}$ of $\mathcal{O}_{i}\setminus\{0\}$, that is
@@
is the semigroup of values of $C_{i}$. Let $\mathcal{O}:=\mathbb{C}\{x,y\}/(f)$. The semigroup of values of $C$ is
@@
Let $C$ be a germ of reduced plane curve at $p\in\mathbb{C}^{2}$. With the previous notation, let $\overline{\varLambda}=\varLambda_{f}\cup\{0\}$. Then, we have